# Short-Context dominance: How Much Local Context Natural Language Actually Needs?

### v1

@@ 6.1 Identifying Long-Context-Relevant Tokens
Validation on NarrativeQA. We validate this hypothesis using NarrativeQA (kocisky2018narrativeqa). We focus on QA pairs where answers are 1–2 words (typically character names or locations), filtering out cases resolvable without the story context. Given a context with the story, question, and partial answer, we classify the ground-truth next token as an Answer Token and all other vocabulary tokens as Non-Answer Tokens. Figure 8 shows LSPS distributions for both categories using LLaMA-3-8B.
Observe the clear separation: Answer tokens exhibit significantly higher LSPS values, confirming that long-context relevant tokens can be identified through their probability shifts. Concretely, using threshold $\epsilon\in[0.05,0.07]$, we capture over 50% of answer tokens while maintaining low false positive rates (<5<5%) on non-answer tokens. Our objective is capturing long-context-relevant tokens while avoiding irrelevant ones, which this approach achieves effectively. In Appx. F, we compare our method to a modified version that uses log-probability ratio (rather than difference) common in prior work (CADBoost; CoherenceBoosting; vanderpoel2022mutual; PerplexityLongCtx), demonstrating the superior stability and precision of our difference-based approach.
@@ 6.2 Targeted Long-Context Token Boosting
Knowing how to (1) detect long-context sequences using LSDS and (2) identify long-context relevant tokens via LSPS, we now combine these tools to improve Q&A performance. The core insight is to promote those identified long-context relevant tokens, downplaying biases from the short context. Simultaneously, this can help downplaying tokens that are assigned high probability as an artifact of noise (sharma2023truth), biases rooted in training data due to potential word/token imbalances (tokenFreqImpact; kassnertokenFreq) and hallucination (hallucination).
@@
Experimental setup. We evaluate on NarrativeQA (kocisky2018narrativeqa), HotpotQA (hotpot), and MultiFieldQA (bai2024longbenchbilingualmultitaskbenchmark) from the LonBench dataset. Focusing on stories with $\geq 1000$ tokens to emphasize long-context dependencies, we sample 3,000 examples from NarrativeQA and HotpotQA while using the full 150 examples from MultiFieldQA. We set $\gamma=0.1225$ [^1] The boosting threshold $\gamma=0.1225$ is intentionally more liberal than the classification threshold $\tau=0.6$ used in Sec. 5. While $\tau$ was set conservatively for clear evaluation of long vs. short context, $\gamma$ captures borderline cases where targeted boosting may still be beneficial. and $\epsilon=0.05$. We test on LLaMA-2-7B, LLaMA-3-8B, Mistral-7B, and Qwen-2-7B. For each question-text pair, we generate 5 answers using nucleus sampling and report both average and best-of-5 F1/BLEU/ROUGE scores. We compare our TaBoo approach against two baselines: (1) vanilla nucleus sampling, and (2) Context Aware Decoding (CAD) (CoherenceBoosting; CADBoost) with $\alpha=0.5$, which applies probability adjustments to all generations and their tokens, unlike our targeted boosting of only long-context relevant tokens in detected long-context sequences.
Results. Table 1 shows TaBoo consistently outperforms vanilla nucleus sampling across all models and datasets. Compared to CAD, TaBoo achieves superior F1 performance on 11 out of 12 dataset-model combinations, losing only on NarrativeQA with LLaMA-3-8B (although not wrt BLEU score). The improvements generalize across model architectures and scale to higher-performing base models like Qwen2-7B. Best-per-example scores also favor TaBoo, with gains over vanilla ranging from 0.5–8.1 F1 points and generally outperforming CAD as well. Additional results in Appx. F.1, F.4 present ablations on hyperparameter selection ($\gamma,\epsilon,\lambda$). Experiments for short summarization on XSUM (XsumPaper) are deferred to Tab. 7 in the appendix. Additionally, in Appx. E.4 we show that the computational overhead of LSDS (and thus TaBoo) is minimal for long sequences.
@@ 7 Outlook
Our work provides a systematic framework for understanding context dependency in language models, with implications for inference efficiency in tasks like QA through principled post-hoc decoding. Specifically, our ability to detect long-context requirements without ground-truth tokens opens opportunities for context-aware generation methods, but also for more targeted evaluation of long-context capabilities and improved training approaches (e.g., following PerplexityLongCtx). Beyond immediate applications, our findings reveal the short-context dominance hypothesis as an inherent property of natural language sequences, with potential broader implications for understanding how language models process and generate text, such as providing alternative motivations for recently proposed logit-adjusted decoding modifications in the hallucination literature CADBoost.
@@ Appendix B Discussion and Related Work
Pursuit of Long Context: Capturing dependencies that extend beyond a few tokens has been a long-standing difficulty in language modeling. Early statistical and neural models either truncated context to short n-grams or attempted to maintain memory through recurrence, but both approaches faced limitations with sparsity or vanishing gradients (chengoodmanNGram; bengio2003neural; SchmidhuberLSTM; cho2014learningphraserepresentationsusing). The Transformer architecture (vaswani2017attention) marked a turning point, with self-attention providing a scalable mechanism for integrating information from across the entire input. Contemporary open-access models such as LLaMA 3 (grattafiori2024llama3herdmodels), Mistral (jiang2023mistral7b), Qwen2 (yang2024qwen2technicalreport), and Gemma (gemmateam2024gemmaopenmodelsbased) support context lengths from 8K to 128K tokens, large enough to encode entire novels in a single pass. To make such extensions feasible, architectural innovations like rotary position encodings (RoPE) (Rope), attention linear biases (ALiBi) (Alibi), and position interpolation (chen2023extendingcontextwindowlarge) enable models to extrapolate beyond their training horizon, while retrieval-augmented designs (borgeaud2022improving; izacard2022atlasfewshot; wang2023augmentinglanguagemodelslongterm) surface or cache relevant information as an alternative to enlarging the raw attention window. Yet, greater architectural capacity does not imply that models make effective use of long-range context at inference time—the focus of our analysis.
@@
Context Utilization: Despite larger context windows, studies show that LMs often do not utilize long-range information. khandelwal2018sharpfuzzy characterize predictions as “sharp nearby, fuzzy far away,” with sensitivity concentrated in the most recent span, while sun2021longcontext demonstrate that only a small fraction of tokens benefit from context beyond the first few thousand tokens. Such works emphasize the inherent recency bias in language models. Complementing these results, liu2023lostmiddlelanguagemodels document a strong positional bias—models attend to evidence at the edges of long prompts while neglecting the middle, a challenge further analyzed by foundinthemiddle, who trace the effect to rotary positional encodings and propose positional modifications to resolve it. To mitigate these limitations, document retrieval approaches have been proposed to surface relevant passages at inference time (borgeaud2022improving; izacard2022atlasfewshot), while recent work emphasizes the role of training data quality in enabling long-context utilization. In particular, checnLDAMlongcontext introduce an attention-based dependency measurement framework (LADM) to identify long documents with strong internal dependencies, showing that selecting such high-quality data for continual pretraining substantially improves long-context performance. Beyond interventions at the system and data level, attribution-based methods such as selfcite directly test the necessity and sufficiency of context spans, offering a finer-grained perspective on how LLMs actually use long inputs. Similarly, we study the utilization of context from a minimal required sub-context viewpoint, showing that majority of next token queries utilize very local information. Additionally, we provide methodology to identify the minimal required sub context and while distinguishing between long-short context queries.
@@
Contrastive Decoding: Beyond architectural and data-level interventions, several ad hoc inference-time methods aim to improve token generation. Early work encouraged generation that depends more strongly on the given context rather than defaulting to frequent or generic outputs. (diversitypromoting2016) propose a mutual-information–based objective for dialogue generation to discourage generic responses, while (brown2020language) address this issue in multiple-choice QA by normalizing candidate likelihoods against unconditional probabilities, thereby encouraging context-dependent answers. CoherenceBoosting introduced coherence boosting, a general inference-time method that reweights the next-token distribution to favor predictions supported by the full long context by explicitly contrasting it against the distribution induced by a shortened context. vanderpoel2022mutual proposed an entropy-aware decoding strategy for summarization, where the scoring function switches to pointwise mutual information between the source document and the next token when model uncertainty is high, thereby discouraging hallucinations and reducing the tendency to select high-frequency but unsupported tokens. Similar to CoherenceBoosting but operating on a logit level, CADBoost introduce Context-Aware Decoding (CAD), an inference-time method designed to reduce hallucinations by explicitly amplifying the difference between a model’s output probabilities with and without the provided context. Much of this line of research grows out of the broader family of contrastive decoding methods (li2023contrastivedecoding; zhao2024contextual; liu2021dexperts) which are designed for ad hoc modification toe next token distribution to improve language generation. For our case, instead of focusing on performance we first provide an objective study of short context bias of language and use our findings to design intuitive and explainable algorithms to to identify long long contexts, relevant tokens and how to combine this knowledge to help improve generation.
@@
n-gram and its short context relevance: Recent work has revisited n-gram models as complements or interpretive tools for neural language models. ngramisback showed that residual learning with a small n-gram LM can regularize neural text generation and reduce hallucination. More recently, liu2025infinigramscalingunboundedngram introduced Infini-gram, which scales n-gram models to trillions of tokens and supports unbounded context length using suffix arrays, enabling both strong next-token prediction and new diagnostic analyses of neural LMs. In parallel, nguyenngramTransfomer argue for understanding Transformers through the lens of n-grams, showing that simple n-gram rulesets can approximate a majority of model predictions (e.g., covering 68–79% of top-1 predictions across benchmarks). Given that n-grams are inherently short-range models—even Infini-gram typically captures at most 32-token dependencies—the fact that they achieve performance comparable to Transformers on standard text suggests a structural bias in language toward short-context sufficiency, which allows such models to achieve moderate success despite their limited horizon. Our observations regarding the overwhelming prevalence of inherently short contexts in natural language can help explain the reason behind the success of such methods which rely on local information using n-grams for generation.
@@
Long Context Evaluation: Much of the evaluation of a model’s contextual understanding has focused on tasks such as question answering, retrieval, and needle-in-the-haystack probing, evaluated on datasets such as NarrativeQA (kocisky2018narrativeqa), TriviaQA (joshi2017trivia), QuALITY (pang2022qualityquestion), and LongBench (bai2024longbenchbilingualmultitaskbenchmark). While these benchmarks test a model’s ability to extract specific information from distant context, they differ from standard language modeling and tend to be highly task-specific. A recent study by PerplexityLongCtx proposes a method for identifying tokens with long-context dependencies and encourages training-time metrics that distinguish such tokens. While their work focuses on a binary classification of long- vs. short-context tokens, we adopt a more fine-grained perspective: treating the language model as a probabilistic oracle and estimating the minimal context required for each next-token prediction in natural text. Additionally, we provide methods for detection of long-context and tokens with long-context relevance without requiring the actual ground truth next token, making our methods applicable during inference.
@@
Decoding Strategies: Given our assumption that language models serve as strong proxies for language understanding, it is important to account for the decoding strategy used during inference. A growing body of research has shown that different sampling methods—such as greedy decoding, top-$k$ sampling (radford2019language; fan-etal-2018-hierarchical), nucleus ($p$) sampling (holtzman2020curious), and adaptive techniques (basu2021mirostatneuraltextdecoding; zhu2024adaptive)—can substantially influence output diversity, factuality, and calibration. Greedy decoding in particular has been shown to produce degenerate or overly deterministic outputs, while adaptive and dynamic approaches aim to adjust sampling entropy and generate a high-quality, contextually valid subset of tokens (holtzman2020curious; zhu2024adaptive; basu2021mirostatneuraltextdecoding). When treating the language model as a statistical oracle for analyzing context usage, it is essential to consider how decoding strategy influences conclusions about effective context length. This perspective may help improve the practical utility of methods such as PerplexityLongCtx, which focus primarily on a single sample from the next-token distribution to classify tokens by their context length requirements. Accordingly, we provide a dedicated analysis of how decoding strategies impact context dependence in next-token prediction and its relation to short-long context detection.
@@ C.2 Language and Domain Results
We conduct two sets of experiments to confirm the robustness of the short-context dominance with respect toe different languages and domains. First, we examine the language effect by running inference on Wikipedia article translations in Arabic, Chinese, French, German, Korean, Russian, and Thai, using the Qwen2-7B model (yang2024qwen2technicalreport). These texts are sampled from the Wikimedia dataset (wikimedia), and follow the same selection and truncation procedure as our English WikiText documents. Results in Table 3 and Fig. 12 confirm that the reliance on short context is preserved across languages, with similarly highly skewed token count distributions and consistent heavy-tailed fits. Additionally, we perform a small test analyzing the Chain-of-Thought (CoT) generated text by the S1 model (S1Reasoningpaper), analyzing the MCL using Qwen-2.5-7B to validate the short-context dominance. Noticeably, looking at the exponential decay in Fig. 11, we see that the CoT contexts exhibit a strong short-context dominance.
@@
Second, to evaluate domain knowledge effects, we test on three specialized datasets: CCDV PubMed Summarization (cohan-etal-2018-discourse) for biomedical abstracts, Open Web Math (paster2023openwebmath)for mathematical content, and LCC_Python (microsoft2024lccpython) for code documents, across all 3 models: LLaMA-3-8B (grattafiori2024llama3herdmodels), Mistral-7B-Instruct (v0.1) (jiang2023mistral7b) and Qwen2-7B (yang2024qwen2technicalreport). More specifically, the contexts selected from the math and medical datasets are within 1024 tokens, while the context length varies from 6000 to 7000 tokens on the code dataset. As shown in Fig. 13 and Fig. 14, the short-context dominance pattern remains intact across all three domains and models. This suggests models’ reliance on local context persist even when dealing with knowledge-intensive content.

### v2

@@ 6.1 Identifying Long-Context-Relevant Tokens
Validation on NarrativeQA. We validate this hypothesis using NarrativeQA (Kočiský et al., 2018). We focus on QA pairs where answers are 1–2 words (typically character names or locations), filtering out cases resolvable without the story context. Given a context with the story, question, and partial answer, we classify the ground-truth next token as an Answer Token and all other vocabulary tokens as Non-Answer Tokens. Figure 8 shows LSPS distributions for both categories using LLaMA-3-8B.
Observe the clear separation: Answer tokens exhibit significantly higher LSPS values, confirming that long-context relevant tokens can be identified through their probability shifts. Concretely, using threshold $\epsilon\in[0.05,0.07]$, we capture over 50% of answer tokens while maintaining low false positive rates (<5<5%) on non-answer tokens. Our objective is capturing long-context-relevant tokens while avoiding irrelevant ones, which this approach achieves effectively. In Appx. F, we compare our method to a modified version that uses log-probability ratio (rather than difference) common in prior work (Duh et al., 2024; Malkin et al., 2022; van der Poel et al., 2022; Fang et al., 2025), demonstrating the superior stability and precision of our difference-based approach.
@@ 6.2 Targeted Long-Context Token Boosting
Knowing how to (1) detect long-context sequences using LSDS and (2) identify long-context relevant tokens via LSPS, we now combine these tools to improve Q&A performance. The core insight is to promote those identified long-context relevant tokens, downplaying biases from the short context. Simultaneously, this can help downplaying tokens that are assigned high probability as an artifact of noise (Sharma et al., 2023), biases rooted in training data due to potential word/token imbalances (Razeghi et al., 2022; Kassner et al., 2020) and hallucination (Ji et al., 2023b; Bender et al., 2021).
@@
Experimental setup. We evaluate on NarrativeQA (Kočiský et al., 2018), HotpotQA (Yang et al., 2018), and MultiFieldQA (Bai et al., 2024) from the LonBench dataset. Focusing on stories with $\geq 1000$ tokens to emphasize long-context dependencies, we sample 3,000 examples from NarrativeQA and HotpotQA while using the full 150 examples from MultiFieldQA. We set $\gamma=0.1225$ [^1] The boosting threshold $\gamma=0.1225$ is intentionally more liberal than the classification threshold $\tau=0.6$ used in Sec. 5. While $\tau$ was set conservatively for clear evaluation of long vs. short context, $\gamma$ captures borderline cases where targeted boosting may still be beneficial. and $\epsilon=0.05$. We test on LLaMA-2-7B, LLaMA-3-8B, Mistral-7B, and Qwen-2-7B. For each question-text pair, we generate 5 answers using nucleus sampling and report both average and best-of-5 F1/BLEU/ROUGE scores. We compare our TaBoo approach against two baselines: (1) vanilla nucleus sampling, and (2) Context Aware Decoding (CAD) (Malkin et al., 2022; Duh et al., 2024) with $\alpha=0.5$, which applies probability adjustments to all generations and their tokens, unlike our targeted boosting of only long-context relevant tokens in detected long-context sequences.
Results. Table 1 shows TaBoo consistently outperforms vanilla nucleus sampling across all models and datasets. Compared to CAD, TaBoo achieves superior F1 performance on 11 out of 12 dataset-model combinations, losing only on NarrativeQA with LLaMA-3-8B (although not wrt BLEU score). The improvements generalize across model architectures and scale to higher-performing base models like Qwen2-7B. Best-per-example scores also favor TaBoo, with gains over vanilla ranging from 0.5–8.1 F1 points and generally outperforming CAD as well. Additional results in Appx. F.1, F.4 present ablations on hyperparameter selection ($\gamma,\epsilon,\lambda$). Experiments for short summarization on XSUM (Narayan et al., 2018) are deferred to Tab. 7 in the appendix. Additionally, in Appx. E.4 we show that the computational overhead of LSDS (and thus TaBoo) is minimal for long sequences.
@@ 7 Outlook
Our work provides a systematic framework for understanding context dependency in language models, with implications for inference efficiency in tasks like QA through principled post-hoc decoding. Specifically, our ability to detect long-context requirements without ground-truth tokens opens opportunities for context-aware generation methods, but also for more targeted evaluation of long-context capabilities and improved training approaches (e.g., following Fang et al. (2025)). Beyond immediate applications, our findings reveal the short-context dominance hypothesis as an inherent property of natural language sequences, with potential broader implications for understanding how language models process and generate text, such as providing alternative motivations for recently proposed logit-adjusted decoding modifications in the hallucination literature Duh et al. (2024).
@@ Appendix B Discussion and Related Work
Pursuit of Long Context: Capturing dependencies that extend beyond a few tokens has been a long-standing difficulty in language modeling. Early statistical and neural models either truncated context to short n-grams or attempted to maintain memory through recurrence, but both approaches faced limitations with sparsity or vanishing gradients (Chen and Goodman, 1996; Bengio et al., 2003; Hochreiter and Schmidhuber, 1997; Cho et al., 2014). The Transformer architecture (Vaswani et al., 2017) marked a turning point, with self-attention providing a scalable mechanism for integrating information from across the entire input. Contemporary open-access models such as LLaMA 3 (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), Qwen2 (Yang et al., 2024), and Gemma (Team et al., 2024) support context lengths from 8K to 128K tokens, large enough to encode entire novels in a single pass. To make such extensions feasible, architectural innovations like rotary position encodings (RoPE) (Su et al., 2021), attention linear biases (ALiBi) (Press et al., 2021), and position interpolation (Chen et al., 2023) enable models to extrapolate beyond their training horizon, while retrieval-augmented designs (Borgeaud et al., 2022; Izacard et al., 2022; Wang et al., 2023) surface or cache relevant information as an alternative to enlarging the raw attention window. Yet, greater architectural capacity does not imply that models make effective use of long-range context at inference time—the focus of our analysis.
@@
Context Utilization: Despite larger context windows, studies show that LMs often do not utilize long-range information. Khandelwal et al. (2018) characterize predictions as “sharp nearby, fuzzy far away,” with sensitivity concentrated in the most recent span, while Sun et al. (2021) demonstrate that only a small fraction of tokens benefit from context beyond the first few thousand tokens. Such works emphasize the inherent recency bias in language models. Complementing these results, Liu et al. (2023) document a strong positional bias—models attend to evidence at the edges of long prompts while neglecting the middle, a challenge further analyzed by Zhang et al. (2024), who trace the effect to rotary positional encodings and propose positional modifications to resolve it. To mitigate these limitations, retrieval-based methods surface relevant passages at inference time (Borgeaud et al., 2022; Izacard et al., 2022), while recent work emphasizes the role of training data quality in enabling long-context utilization. In particular, Chen et al. (2025) introduce an attention-based dependency measurement framework (LADM) to identify long documents with strong internal dependencies, showing that selecting such high-quality data for continual pretraining substantially improves long-context performance. Beyond interventions at the system and data level, attribution-based methods such as Chuang et al. (2025) directly test the necessity and sufficiency of context spans, offering a finer-grained perspective on how LLMs actually use long inputs. Similarly, we study the utilization of context from a minimal required sub-context viewpoint, showing that majority of next token queries utilize very local information. Additionally, we provide methodology to identify the minimal required sub context, based on ground truth next token or next token distribution matching, rather than only distinguishing and separating long-short context queries.
@@
Contrastive Decoding: Beyond architectural and data-level interventions, several ad hoc inference-time methods aim to improve token generation. Early work encouraged generation that depends more strongly on the given context rather than defaulting to frequent or generic outputs. (Li et al., 2016) propose a mutual-information–based objective for dialogue generation to discourage generic responses, while (Brown et al., 2020) address this issue in multiple-choice QA by normalizing candidate likelihoods against unconditional probabilities, thereby encouraging context-dependent answers. Malkin et al. (2022) introduced coherence boosting, a general inference-time method that reweights the next-token distribution to favor predictions supported by the full long context by explicitly contrasting it against the distribution induced by a shortened context. van der Poel et al. (2022) proposed an entropy-aware decoding strategy for summarization which revolves around promoting tokens that appear more often in the source text whenever next-token-entropy (model uncertainty) is high, thereby discouraging hallucinations and reducing the tendency to select high-frequency tokens based on the model’s own biases (promoting high-frequency tokens in language which don’t have a strong correlation to the source material). Similar to Malkin et al. (2022) but operating on a logit level, Duh et al. (2024) introduce Context-Aware Decoding (CAD), an inference-time method designed to reduce hallucinations by explicitly amplifying the difference between a model’s output probabilities with and without the provided context. Much of this line of research grows out of the broader family of contrastive decoding methods (Li et al., 2023; Zhao et al., 2024; Liu et al., 2021) which are designed for ad hoc modification toe next token distribution to improve language generation. For our case, instead of focusing on performance we first provide an objective study of short context bias of language and use our findings to design intuitive and explainable algorithms to to identify long long contexts, relevant tokens and how to combine this knowledge to help improve generation.
@@
n-gram and its short context relevance: Recent work has revisited n-gram models as complements or interpretive tools for neural language models. Li et al. (2022) showed that residual learning with a small n-gram LM can regularize neural text generation and reduce hallucination. More recently, Liu et al. (2025) introduced Infini-gram, which scales n-gram models to trillions of tokens and supports unbounded context length using suffix arrays, enabling both strong next-token prediction and new diagnostic analyses of neural LMs. In parallel, Nguyen (2024) argue for understanding Transformers through the lens of n-grams, showing that simple n-gram rulesets can approximate a majority of model predictions (e.g., covering 68–79% of top-1 predictions across benchmarks). Given that n-grams are inherently short-range models—even Infini-gram typically captures at most 32-token dependencies—the fact that they achieve performance comparable to Transformers on standard text suggests a structural bias in language toward short-context sufficiency, which allows such models to achieve moderate success despite their limited horizon. Our observations regarding the overwhelming prevalence of inherently short contexts in natural language can help explain the reason behind the success of such methods which rely on local information using n-grams for generation.
@@
Long Context Evaluation: Much of the evaluation of a model’s contextual understanding has focused on tasks such as question answering, retrieval, and needle-in-the-haystack probing, evaluated on datasets such as NarrativeQA (Kočiský et al., 2018), TriviaQA (Joshi et al., 2017), QuALITY (Pang et al., 2022), and LongBench (Bai et al., 2024). While these benchmarks test a model’s ability to extract specific information from distant context, they differ from standard language modeling and tend to be highly task-specific. A recent study by Fang et al. (2025) proposes a method for identifying tokens with long-context dependencies and encourages training-time metrics that distinguish such tokens. Following a similar pattern, Helm et al. (2025) uses a similar strategy to promote long context tokens during pretraining. While such work focuses on a binary classification of long- vs. short-context tokens and often consider windows of size 8,000 tokens as the "short-context" (compared to our focus on smaller windows of 32 tokens), we adopt a more fine-grained perspective: treating the language model as a probabilistic oracle and estimating the minimal context required for each next-token prediction in natural text. Additionally, we provide methods for detection of long-context and tokens with long-context relevance without requiring the actual ground truth next token, making our methods applicable during inference. While such works focus more on the long context retrieval capabilities of models and its practical implications, we analyze the short context bias of language itself, illustrating the short-context dominance of human language itself and provide one example of how to combat such biases using TaBoo.
@@
Decoding Strategies: Given our assumption that language models serve as strong proxies for language understanding, it is important to account for the decoding strategy used during inference. A growing body of research has shown that different sampling methods—such as greedy decoding, top-$k$ sampling (Radford et al., 2019; Fan et al., 2018), nucleus ($p$) sampling (Holtzman et al., 2020), and adaptive techniques (Basu et al., 2021; Zhu et al., 2024)—can substantially influence output diversity, factuality, and calibration. Greedy decoding in particular has been shown to produce degenerate or overly deterministic outputs, while adaptive and dynamic approaches aim to adjust sampling entropy and generate a high-quality, contextually valid subset of tokens (Holtzman et al., 2020; Zhu et al., 2024; Basu et al., 2021). When treating the language model as a statistical oracle for analyzing context usage, it is essential to consider how decoding strategy influences conclusions about effective context length. This perspective may help improve the practical utility of methods such as Fang et al. (2025) and Helm et al. (2025), which focus primarily on a single sample from the next-token distribution to classify tokens by their context length requirements. Accordingly, we provide a dedicated analysis of how decoding strategies impact context dependence in next-token prediction and its relation to short-long context detection.
@@ C.2 Language and Domain Results
We conduct two sets of experiments to confirm the robustness of the short-context dominance with respect toe different languages and domains. First, we examine the language effect by running inference on Wikipedia article translations in Arabic, Chinese, French, German, Korean, Russian, and Thai, using the Qwen2-7B model (Yang et al., 2024). These texts are sampled from the Wikimedia dataset (Schwenk et al., 2019), and follow the same selection and truncation procedure as our English WikiText documents. Results in Table 3 and Fig. 12 confirm that the reliance on short context is preserved across languages, with similarly highly skewed token count distributions and consistent heavy-tailed fits. Additionally, we perform a small test analyzing the Chain-of-Thought (CoT) generated text by the S1 model (Muennighoff et al., 2025), analyzing the MCL using Qwen-2.5-7B to validate the short-context dominance. Noticeably, looking at the exponential decay in Fig. 11, we see that the CoT contexts exhibit a strong short-context dominance.
@@
Second, to evaluate domain knowledge effects, we test on three specialized datasets: CCDV PubMed Summarization (Cohan et al., 2018) for biomedical abstracts, Open Web Math (Paster et al., 2023)for mathematical content, and LCC_Python (Microsoft, 2024) for code documents, across all 3 models: LLaMA-3-8B (Grattafiori et al., 2024), Mistral-7B-Instruct (v0.1) (Jiang et al., 2023) and Qwen2-7B (Yang et al., 2024). More specifically, the contexts selected from the math and medical datasets are within 1024 tokens, while the context length varies from 6000 to 7000 tokens on the code dataset. As shown in Fig. 13 and Fig. 14, the short-context dominance pattern remains intact across all three domains and models. This suggests models’ reliance on local context persist even when dealing with knowledge-intensive content.